OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form. A. y = -2x + 14 B. y = -2x + 5 C. y = -2x + 9 D. y = -2x + 6

OpenStudy (haseeb96):

A is the correct answer y-4=-2x+10 y=-2x+10+4 y=-2x+14

OpenStudy (anonymous):

@Haseeb96 Yay! Thank yoU!!!